### v3

# WebDRO: A Web-based Group-level Clustering and Reweighting Method for Unsupervised Dense Retrieval

Abstract
The anchor-document data derived from web graphs offers a wealth of paired information for training dense retrieval models in an unsupervised manner. However, the presence of inherent noise invariably compromises the robustness of training dense retrieval models, consequently hurting the performance. In this paper, we introduce WebDRO, an efficient approach for clustering the web graph data and optimizing group weights to enhance the robustness of the pretraining process of dense retrieval models on web graphs. Initially, we build an embedding model for clustering anchor-document pairs. Specifically, we contrastively train the embedding model for link prediction, which guides the embedding model in capturing the inherent document features behind the web graph links. Subsequently, we employ the group distributional robust optimization to recalibrate the weights across different clusters of anchor-document pairs during training dense retrieval models, directing the model to assign higher weights to clusters with higher loss and focus more on worst-case scenarios. Our experiments conducted on MS MARCO and BEIR demonstrate that our method can effectively improve retrieval performance in unsupervised training settings. Further analysis confirms the stability and validity of group weights learned by WebDRO. All codes will be released via GitHub.
WebDRO: A Web-based Group-level Clustering and Reweighting Method for Unsupervised Dense Retrieval
Peixuan Hanthanks: Part of this work was done while the author was visiting Carnegie Mellon University. Tsinghua University hanpx20@mails.tsinghua.edu.cn                        Zhenghao Liu Northeastern University liuzhenghao@mail.neu.edu.cn
Zhiyuan Liu Tsinghua University liuzy@tsinghua.edu.cn                        Chenyan Xiong Carnegie Mellon University cx@cs.cmu.edu
1 Introduction
Dense retrieval models, which encode the semantics of text into dense vectors, have become prevalent and achieved state-of-the-art performance on many public benchmarks thanks to the surge of pre-trained language models like BERT Devlin et al. (2019) and T5 Raffel et al. (2020). With enough annotated data, dense retrieval models can adeptly capture semantic features in supervised scenarios, outperforming traditional sparse methods such as BM25 Robertson et al. (2009, 1995), which solely retrieves documents containing keywords already present in the query. However, annotated data is inaccessible in many scenarios either due to the extensive effort needed for annotation or privacy concerns. This highlights the necessity for unsupervised dense retrievers. Several paradigms of unsupervised dense retrieval have been proposed, among which leveraging web anchors is a promising choice. Specifically, anchor-document pairs can approximate supervised signals where the document contains relevant information that appears in the query better than most existing methods Xie et al. (2023a); Zhang et al. (2020b).
Web anchors contain both useful signals and noise, limiting the effectiveness of anchor-based training methods Zhang et al. (2020b). Previous research has primarily focused on instance-level selection to distinguish useful contrastive pairs from uninformative ones by assigning a weight to each instance Moore and Lewis (2010); Zhang et al. (2020a). Nevertheless, directly weighting at the instance level is often coarse, since the information contained in individual data points is insufficient. To reach a more tailored weighting, some approaches leverage feedback signals during the training process, such as Active Learning Settles (2011); Coleman et al. (2020) and Meta Adaptive Ranking Sun et al. (2021). However, these methods have the drawback of requiring the model to be trained multiple times, which is highly time-consuming and computationally intensive.
We address the lack of both efficient and accurate retrieval data selection methods by introducing a group-level clustering and reweighting strategy–WebDRO. WebDRO enables us to conduct data selection with the feedback signal (training loss) in a single pass, as well as, cluster data into groups due to that the group encapsulates significantly more information than a single data point. In other words, we are “smoothing” the weights within each group through clustering, thereby enhancing the quality and stability of reweighting. Specifically, we employ a two-step approach to train an unsupervised dense retrieval model using WebDRO. In the first step, we utilize structural information from the web graph to train an embedding model with contrastive learning, which plays a role as the link predictor. By leveraging the embedding model to conduct vector representations for documents in the anchor-document pairs, we can produce high-quality clusters of contrastive pairs. In the second step, we utilize Group Distributionally Robust Optimization (GroupDRO) Oren et al. (2019); Sagawa et al. (2019) to reweight the clusters. Originally designed to tackle distributional shifts in language model training, GroupDRO dynamically adjusts the weights of different groups based on the difficulty of learning with that group. Consequently, the model can effectively discern informative web pages from noisy, uninformative ones.
We conduct experiments on two widely used retrieval datasets, MS MARCO Nguyen et al. (2016) and BEIR Thakur et al. (2021). Results indicate that our method notably enhances retrievers’ performance in both unsupervised and finetuned settings. Specifically, WebDRO yields a performance improvement of 1.2% (measured by nDCG@10) on MS MARCO and 1.1% on BEIR compared to our main baseline, Anchor-DR Xie et al. (2023b), in an unsupervised setting. Analytical experiments further demonstrate that our model consistently produces stable, high-quality clusters and accurately identifies groups of high importance. Additionally, our method exhibits notable efficiency and scalability advantages, rendering it particularly well-suited for pretraining based on large-scale corpus.
2 Related Work
This section describes relevant work, focusing on dense retrieval models, data selection methods, and distributionally robust optimization.
Dense Retrieval. Dense Retrieval is an approach where both the query and document are separately encoded through a neural network. The prevalent training approach, introduced by Karpukhin et al. (2020), trains the network with a contrastive loss Although dense retrieval has over-performed sparse methods Robertson et al. (2009) in many tasks, its benefits are generally confined to well-annotated data. Most unsupervised dense retrieval methods focus on using unannotated data to generate query-document pairs that semantically resemble contrastive pairs in supervised scenarios, like ICT Izacard et al. (2021) and COCO-LM Meng et al. (2021). Hyperlinks on the Internet contain rich information about relationships between texts, and thus can also be utilized in dense retrieval. Anchor-DR Xie et al. (2023b) contrastively trains dense retrievers using web anchors-document pairs, which has shown superiority over other unsupervised dense retrieval methods.
Retrieval Data Selection. Most previous work selects retrieval data by comparing the query’s similarity with texts that are considered high-quality. Typically, they assign an attribute to each instance and select a subset based on the attribute’s value Moore and Lewis (2010); Zhang et al. (2020a). However, being similar in form to high-quality text alone does not necessarily make a good query Zhang et al. (2020b). To tailor the selection, some approaches use feedback signals from training as the criterion for selection, namely, Meta Adaptive Ranking Sun et al. (2021) and Active Learning Coleman et al. (2020). The main drawback of using feedback signals is time and computational power consumption since this type of method needs to train the model several times. Such methods are not suitable for pretraining dense retrievers using large-scale datasets.
(a) Web-based clustering.
(b) Contrastive training with WebDRO.
Figure 1: Training process of WebDRO. Figure 0(a) shows the clustering process. We use web links to train an embedding model and use the model to cluster documents. Figure 0(b) shows the training process of the retrieval model. We aim to align the anchor with the corresponding document while pushing apart unrelated ones. Each group is assigned a weight that is dynamically updated.
Distributionally Robust Optimization. To address the issue of distribution shift Agarwal et al. (2021); Wiles et al. (2022), researchers have introduced Distributionally Robust Optimization (DRO)  Ben-Tal et al. (2013); Rahimian and Mehrotra (2019); Lin et al. (2022). The goal of DRO is to optimize the worst-case performance of a model across various data distributions, ensuring its robustness against uncertain distributional changes. Specifically, GroupDRO considers robustness over groups and calculates the worst-case loss over groups by optimizing group weights and model weights concurrently. GroupDRO can be used to train more robust models, as well as to learn the importance of different data groups Xie et al. (2023a). Prior applications of GroupDRO used explicitly pre-partitioned groups, and the number of groups is relatively small ($<50$). In this work, we extend the applicable scope of GroupDRO to a large number of groups clustered by implicit information.
3 Methodology
In this section, we first describe the preliminary of dense retrieval (Sec. 3.1). Then we introduce our WebDRO method, which is shown in Figure 1. To begin with, we use the method in Sec. 3.2 to train an embedding model using the web links. Then, we cluster the anchor-document pairs based on vector representations produced by the embedding model. Finally, we train a robust dense retrieval model with Group Distributionally Robust optimization (GroupDRO) as introduced in Sec. 3.3.
3.1 Preliminary of Dense Retrieval
Given a query $q$, the dense retrieval task aims to search relevant document $p$ to satisfy the information needs of users.
In dense retrieval, we train a model to encode both queries and documents as dense vectors and then conduct relevance modeling in an embedding space. Through contrastive learning Oord et al. (2018); Chen et al. (2020); He et al. (2020), the encoder is optimized to align the queries with related documents and pull away the irrelevant documents from queries. We use cosine similarity to measure the relevance between queries and documents and then we minimize the following loss function:
$l_{\text{dr}}(q)=-\log\frac{e^{cos(q,p^{+})/\tau}}{e^{cos(q,p^{+})/\tau}+\sum_{i=1}^{k}{e^{cos(q,p^{-}_{i})/\tau}}},$
where $\tau$ is the temperature hyperparameter. $p^{+}$ stands for the related document for query and $p^{-}_{1},...,p^{-}_{k}$ are negative (unrelated) documents.
3.2 Clustering with the Embedding Model
This subsection introduces our web graph-based anchor-document cluster method. It learns the document representations using the document links in the web graph and then clusters the documents according to their semantic features.
Document Relations in Web Graph. The web graph models the document relationship using hyperlinks, which allows users to navigate from one web page to another or within the same web page. It is evident that the documents linked with a hyperlink will likely share similar or related information. As shown in Figure 0(a), these four documents are related and are linked with the anchor hyperlink, e.g. the link between document A and document B indicate that “horses and gorillas have some similarities”, showing an analogy relationship; the links between document B and document C indicate the definitions of “horses” and “herbivorous”, showing the hypernymy relationship.
Embedding Model Training. We train an embedding model using the document relations in the web graph. We encode the documents using pretrained language models and contrastively train the encoder, which is the same as Sec. 3.1. A web page can link to more than one page, so we enumerate each link to construct contrastive pairs as training examples. For each link, we regard the source web page as the query, the target web page as the positive document, and the unlinked documents as negative ones. As a result, the embedding model is trained as a link predictor, which can distinguish linked web pages from non-linked ones.
Document Clustering. We use the embedding model to encode all documents as dense vectors, and then use Mini-Batch K-Means Sculley (2010), a variant of K-Means Krishna and Murty (1999) to cluster embedding vectors of documents into $n$ groups. Some documents have more than one anchor that links to it, these anchor-document pairs are naturally clustered into the same group. To make sure each group has statistical significance, we set a threshold of the number of documents of different groups as $MinSize=128$. All the groups whose size is smaller than $MinSize$ will be aggregated together and considered together as a larger group. As this large group contains heterogeneous elements with different topics, we should not assign a single weight to this group. Therefore, the large group will not be considered in the reweighting, which is discussed in Sec. 3.3.
3.3 WebDRO: Robustly Unsupervised Retrieval Model Training
The anchor-document pairs contain different types of documents. The noises in these pairs Zhang et al. (2020b) may affect the performance of dense retrievers and conduct unstable training process. To alleviate this problem, we follow Sagawa et al. (2019) to reweight different anchor-document groups and robustly train dense retrievers.
Following GroupDRO, each group is assigned with a weight that indicates its importance, and the weights are dynamically updated during the training process. Specifically, we divide the anchor-document pairs into $n$ groups after clustering. Then, for a training instance $p$ that belongs to group $k$, the training loss of WebDRO is calculated as:
$L_{\text{WebDRO}}(p)=l_{\text{dr}}(p)\times w_{k}\times n\times C_{k},$ (2)
where $w_{k}$ is the weight of the $k$-th group, $C_{k}$ is size factor of the $k$-th group, and $l_{\text{dr}}$ is the contrastive loss calculated using Eq. 1. In Figure 0(b), we illustrate 3 clusters, each of them containing documents about the same topic. During contrastive training, group weights, which are multiplied onto the training loss, get dynamically updated.
Then, we describe $C_{k}$ and $w_{k}$, respectively.
Size Factor. In a standard GroupDRO scenario, data are sampled from each group according to the uniform distribution. However, the size of different anchor-document groups in our method may vary significantly. To accommodate this circumstance, WebDRO introduces a size factor $C_{k}$:
$C_{k}=\frac{\sum_{j=1}^{n}{N_{j}}}{n\times N_{k}},$ (3)
where $N_{k}$ is the size of the $k$-th anchor-document group. Smaller groups will have a larger $C_{k}$ value, which serves as a balance for different group sizes.
Group Weighting. We denote the weight of group $k$ at step $t$ as $w_{k}^{t}$. The initial weight of all groups is $w_{k}^{0}=\frac{1}{n}$ and after each step, the sum of group weights should be $1$. In each step, the groups’ weights are updated before the calculation of $L_{\text{WebDRO}}$ (Eq. 2). First, we increase group weights based on training loss. For each group $k$,
$w_{k}^{\prime}=w_{k}^{t-1}\times e^{lr\times C_{k}\times L_{\text{avg}k}}.$ (4)
In this formula, $lr$ is the learning rate of weight updating. $L_{\text{avg}k}$ portraits the contribution of group $k$ in step $t$’s training loss, which can be calculated:
$L_{\text{avg}k}=\frac{\sum_{p\text{ in group }k}{l_{\text{dr}}(p)}}{N_{t}},$ (5)
where $N_{t}$ is the number of training data in step $t$. If all training data in a step belongs to a single group, $L_{\text{avg}k}$ for that group will be the average loss, and other group weights will remain unmodified.
Then, we normalize the modified weights and calculate the weight at step $t$. For each group $k$,
$w_{k}^{t}=\frac{w_{k}^{\prime}}{\sum_{j=1}^{n}{w_{j}^{\prime}}}.$ (6)
During training, we aim to increase the weights of groups with higher contrastive loss, which leads the model to draw more attention to groups that are difficult to learn or contain more information.
Model BM25 Unsupervised Dense Retrievers Finetuned Dense Retrievers
SPAR AugTriever coCondenser Contriever $*{}^{*}$ WebDRO ${}^{\sharp}$ ${}^{\flat}$ WebDRO
MS MARCO 22.8 19.2 20.6 16.2 20.6 26.15 $*{}^{*}$ 36.47 ${}^{\sharp}$ ${}^{\sharp}\flat$
trec-covid 65.6 53.1 53.5 40.4 27.4 72.16 69.26 72.43 73.98 ${}^{\sharp}\flat$
nfcorpus 32.5 26.4 30.3 28.9 31.7 30.70 30.74 29.22 29.75 ${}^{\sharp}\flat$
fiqa 23.6 18.5 22.3 25.1 24.5 23.79 $*{}^{*}$ 26.72 ${}^{\sharp}$ ${}^{\sharp}$
arguana 31.5 42.0 39.1 44.4 37.9 28.39 $*{}^{*}$ 47.59 ${}^{\sharp}$ 48.02
touche2020 36.7 26.1 21.6 11.7 19.3 21.85 $*{}^{*}$ 23.16 24.06 ${}^{\sharp}\flat$
quora 78.9 70.4 82.7 82.1 83.5 85.57 83.36 85.28 85.48 ${}^{\sharp}$
scidocs 15.8 13.4 14.7 13.6 14.9 13.42 $*{}^{*}$ 14.95 15.41 15.30
scifact 66.5 62.6 64.4 56.1 64.9 58.84 $*{}^{*}$ 57.54 ${}^{\sharp}$ ${}^{\sharp}$
nq 32.9 26.2 27.2 17.8 25.4 28.83 $*{}^{*}$ 42.48 ${}^{\sharp}$ ${}^{\sharp}\flat$
hotpotqa 60.3 57.2 47.9 34.0 48.1 53.81 52.27 50.25 ${}^{\sharp}$ ${}^{\sharp}$
dbpedia-entity 31.3 28.1 29.0 21.5 29.2 34.61 $*{}^{*}$ 34.01 ${}^{\sharp}$ ${}^{\sharp}\flat$
fever 75.3 56.9 59.7 61.5 68.2 72.15 $*{}^{*}$ 68.42 ${}^{\sharp}$ 70.93
climate-fever 21.4 16.4 17.7 16.9 15.5 18.87 21.05 18.45 18.95 18.89
cqadupstack 29.9 27.9 27.1 30.9 28.4 28.82 $*{}^{*}$ 33.64 ${}^{\sharp}$ ${}^{\sharp}$
trec-news 39.8 - - 25.4 35.2 41.18 40.24 39.68 ${}^{\sharp}$ ${}^{\sharp}$
signal1m 33.0 - - 21.4 25.0 22.99 $*{}^{*}$ 26.48 27.56 27.82
bioasq 46.5 - - 22.7 32.7 33.88 $*{}^{*}$ 30.07 33.73 ${}^{\sharp}\flat$
robust04 40.8 - - 29.8 32.7 41.01 37.60 ${}^{\flat}$ 43.23 ${}^{\flat}$
Avg. on BEIR14111Average score of 14 public datasets in BEIR. 43.0 36.2 37.0 33.4 36.0 40.84 42.15 43.15 45.14 46.03
Avg. All 41.3 - - 31.6 35.0 38.80 39.91 41.11 43.09 43.89
Table 1: Retrieval results of different models on MS MARCO and BEIR. All models are trained in the unsupervised setting and are evaluated using nDCG@10. The best result on each task is marked in bold, and the best result among unsupervised dense retrievers is underlined222 * indicate statistically significant improvements over Anchor-DR in unsupervised setting. $\sharp$ and $\flat$ indicate that over T5-base and Anchor-DR respectively in finetuned setting..
4 Experimental Methodology
This section introduces the training dataset, evaluation, baseline models, and implementation details.
Training and Evaluation Datasets. In our experiments, we use the anchor document pairs from ClueWeb22 Overwijk et al. (2022), a dataset containing extensive web page information, to train dense retrievers in an unsupervised setting. We use MS MARCO Nguyen et al. (2016) and BEIR Thakur et al. (2021), two widely used English retrieval benchmarks, to evaluate the retrieval effectiveness in both unsupervised and finetuned settings. Following previous work Thakur et al. (2021); Xie et al. (2023b), nDCG@10 is used as our evaluation metric.
Implementation Details. Our implementation is based on OpenMatch Liu et al. (2021); Yu et al. (2023). During training the embedding model for clustering, we consider that the URL of web pages contains valuable information to learn the relations among web pages. Therefore, we incorporate the URL with the title and content as the text representation of web pages. We cluster training data into $500$ groups, yet the actual data may contain slightly less than 500 groups because the groups smaller than $128$ are merged. An exception is DRO w/ URL, which contains $4,732$ groups.
For retrieval models, we start from the T5-base Raffel et al. (2020) ($220$M parameters) checkpoint and contrastively train the dense retrieval model on web data with in-batch negatives and BM25 retrieved negatives for 1 epoch. Then we reinitialize the group weight and continuously train the model using hard (more challenging) negatives Xiong et al. (2021) for 1 more epoch. To enhance the stability of WebDRO, we use gradient accumulation and update the group weights every $500$ training step (gradients of model parameters are not cached).
Baseline Models. We compare WebDRO with a commonly used sparse retriever, BM25 Robertson et al. (2009), and five unsupervised dense retrieval models. Anchor-DR Xie et al. (2023b) regards anchor texts as queries and uses anchor-document pairs without clustering. We describe the other four models in Appendix A.2. We implement two variants of our method by using different clustering strategies. DRO w/ URL directly clusters the web pages that share the same domain name into the same group. DRO w/ Anchor-DR uses Anchor-DR to encode the documents for clustering instead of using a task-specifically trained embedding model. These variants are discussed in Sec. 5.2.
We also conduct experiments to assess WebDRO in a fine-tuned setting. We compare three different base models: T5-base, Anchor-DR, and WebDRO. These models are fine-tuned using the training set from MS MARCO Passage for 1 epoch.
5 Evaluation Results
In this section, we first show the performance of different retrieval models and analyze different clustering methods in our approach. Then we analyze the validity of the group weights learned by WebDRO and compare WebDRO’s performance in different tasks. Finally, some case studies are shown.
5.1 Overall Performance
Table 2 shows the unsupervised retrieval results of WebDRO on MS MARCO and BEIR.
Overall, WebDRO outperforms other dense retrieval models by achieving at least a 1.2% improvement on MS MARCO and a 1.1% improvement on the average of BEIR. By utilizing group weights, WebDRO significantly enhances retrieval performance on most datasets, indicating its effectiveness in filtering out noise in anchor-document pairs and pretraining dense retrievers in a robust and efficient way.
As shown in the right part of Table 2, WebDRO significantly outperforms Anchor-DR and T5-base on most datasets when they are all finetuned with MS MARCO data. Additionally, finetuned WebDRO also outperforms BM25 on 10 sub-tasks (of 19) and has an average performance gain of 2.6%. The performance gain of WebDRO further indicates the advantages of our clustering and reweighting method can be brought to the fine-tuned scenarios.
We conduct our experiment five times on WebDRO using different random seeds and report the average score in Table 2. The disparity between the top-performing and lowest-performing models on MS MARCO is less than $0.3\%$. We compute the cosine similarity between the group weight vectors of five WebDRO models and the lowest similarity score observed is $98.968\%$. This illustrates WebDRO’s ability to learn stable group weights, thereby ensuring robust performance and improving generalizability across diverse datasets.
Model MS MARCO Avg. on BEIR
Anchor-DR 26.15 39.50
DRO w/ URL 26.97 39.79
DRO w/ Anchor-DR 26.64 39.65
WebDRO 27.37 40.60
Table 2: Performance of variants of WebDRO using different clustering strategies in an unsupervised setting.
5.2 Analyses on Clustering Strategy
In the experiment, we conduct experiments to compare the effectiveness of different document clustering strategies, as shown in Table 2.
WebDRO leverages the extensive document relations within the web graph to train an embedding model for clustering. However, DRO w/ Anchor-DR directly employs a dense retrieval model (Anchor-DR) for clustering. As shown in the evaluation results, WebDRO outperforms DRO w/ Anchor-DR by achieving about a 1% improvement. We hypothesize that this disparity arises from the fact that the clustering task concerns the relations between documents, whereas Anchor-DR is trained with anchor-document pairs, emphasizing the modeling of relations between queries and documents. A link predictor style embedding model focuses more on extracting document features to model the relation between web pages, making it more tailored for clustering documents.
URLs are commonly employed to encapsulate the content of a web page, thus DRO w/ URL exhibits an improvement over vanilla Anchor-DR. However, in comparison to DRO w/ URL, WebDRO shows better retrieval performance, illustrating that titles and contents also play a crucial role in enhancing the URL-based representations of web pages. Another reason for DRO w/ URL’s performance decline compared to WebDRO is its coarse clustering for the documents belong to large domains. For instance, all web pages within Wikipedia will be clustered in one single group (this group contains roughly $1$M anchors, constituting $7.2\%$ of whole training data) regardless of their topics.
Figure 2: The relationship between performance gain of WebDRO and performance of Anchor-DR. Each point represents a dataset in BEIR. Datasets with high or negative gains are annotated.
5.3 Retrieval Effectiveness in Different Testing Scenarios
This section focuses on evaluating the role of the group weighting mechanism in training dense retrieval models. In Figure 2, we present the performance points of all 18 datasets from BEIR and show the relationship between Anchor-DR’s performance and the relative performance enhancement achieved by WebDRO over Anchor-DR.
The evaluation results show that the performance gap between Anchor-DR and WebDRO tends to be larger when the baseline nDCG@10 is low. It illustrates that WebDRO performs particularly well on difficult retrieval tasks, like scidocs (11.26%), hotpotQA (19.60%) and cqadupstack (14.39%). Conversely, on datasets where the baseline already exhibits strong performance, such as trec-covid (-4.30%) and quora (-2.58%), the model’s enhancements are marginal or even negative. This aligns with the optimization goal of WebDRO, which is that the groups with larger training losses are assigned larger weights to improve the overall performance across all datasets. This phenomenon suggests that WebDRO’s unique optimization strategy effectively identifies and prioritizes challenging tasks during pretraining, offering a strategic advantage for performance gains, particularly in scenarios where traditional models struggle.
5.4 Effectiveness of the Group Weights Learned by WebDRO
In this section, two experiments are conducted to show the effectiveness of the group weights that are learned by WebDRO.
(a) The weight changes of
50 random groups.
(b) The weight changes of top and bottom groups.
Figure 3: The visualization of group weights during training. Group weights are recorded every 500 steps.
Trend and Distribution of Group Weights. As shown in Figure 3, we explore the distributions of group weights that learned by WebDRO.
We first randomly sample 50 groups to track the changes of group weights during training. As seen in Figure 2(a), the color gradient within each group gradually shifts towards darker, indicating an approximate monotonic change in group weights. Then we rank the groups according to the weights that are assigned by our final WebDRO model and select the top-ranked and bottom-ranked 30 groups respectively to observe the trend of weight changes during training. As shown in Figure 2(b), the weights of top-weighted groups and bottom-weighted groups are consistently increased and decreased, further showing the stability of the group weights learned by WebDRO.
Effectiveness of Retrieval Models Trained with Different Weighted Groups. We conduct experiments to pretrain dense retrieval models using groups assigned with different weights derived from WebDRO. Specifically, we select top and bottom 30 groups based on their weight ranks, and then sample 500k anchor-document pairs from top and bottom groups respectively. Subsequently, we utilize these subsets of training data to train dense retrieval models. For comparison, we also train a model using a randomly sampled set of 500k pairs.
Training Data MS MARCO Avg. on BEIR
All Groups 27.37 40.60
Top Groups 24.12 38.53
Bottom Groups 20.70 34.70
Random 23.42 36.99
Table 3: Performance of retrieval model pretrained with different groups. Models are evaluated with nDCG@10.
As shown in Table 3, the retrieval model trained with top-weighted groups significantly outperforms other models trained with an equal number of 500k anchor-document pairs. It illustrates that these groups that are assigned with high weights contain valuable training signals for training the dense retrievers. The retrieval performance of the model that is trained with randomly sampled training data significantly decreases when we replace the training data with instances in bottom-weighted groups. This highlights the capacity of WebDRO to effectively filter out noise present in the training data.
(a) Top-Weighted Anchors
Anchor: The World Cup Document: In all, 80 nations have played in at least one World Cup. Of these, eight national teams have won the World Cup, and they have added stars to their badges, with each star representing a World Cup victory, with only one exception…
Anchor: Sepolture Premature Document: Have People Been Buried Alive? And could it still happen? Snopes Staff Published 9 June 1999 Image via Shutterstock 10 Claim People have been buried alive by mistake…
(b) Bottom-Weighted Anchors
Anchor: Blood Document: What is an inferior vena cava filter placement? An inferior vena cava filter is a small device that can stop blood clots from traveling through the vena cava vein to the heart and lungs…
Anchor: Evidence Based Physical Therapy Document: Evidence-based physical therapy (EBPT) has been defined as physiotherapy informed by relevant high quality clinical research. The practice of evidence-based physiotherapy should be informed by the integration of…
Table 4: Examples of anchor-document pairs in top-weighted 30 groups and bottom-weighted 30 groups.
5.5 Analyses of the Loss Landscapes
This experiment shows the WebDRO’s superiority from the perspective of the loss landscape. Loss landscape is a common tool for assessing a model’s stability and generalization capabilities Li et al. (2018); Hao et al. (2019). Specifically, we apply small perturbations to the model’s parameters from two random directions and visualize the losses after perturbations in a plane. The “flatness” of the loss landscape is defined as the size of the region around the origin where the training loss remains low Hochreiter et al. (1997). A flatter loss landscape usually indicates better robustness to noises and a higher chance of converging to the optimal when randomly initialized.
As shown in Figure 4, it is evident that with increasing perturbation, the training loss in WebDRO exhibits a gradual increase, whereas in Anchor-DR, the rise in training loss is steep in specific directions. Besides, it can be observed that WebDRO encompasses a larger area with low training loss (the area inside the red lines) than Anchor-DR. In summary, WebDRO presents a flatter and more uniform landscape than Anchor-DR, indicating enhanced stability and efficiency in training and heightened generalization capabilities.
(a) Anchor-DR.
(b) WebDRO.
Figure 4: Loss landscapes of Anchor-DR and WebDRO. We constrain the perturbation in each direction to $(-0.3,0.3)$. Each closed curve in the figure indicates that the corresponding positions have the same loss.
5.6 Case Study
Finally, we show some case studies in Table 4.
We sample examples from both the top-weighted and bottom-weighted groups. In the top-weighted groups, the majority of anchor-document pairs exhibit the following characteristics: both the anchor and the document contain rich information, and the anchor text serves as the primary topic or theme of the document. A large proportion of anchor-documents pairs in bottom groups exhibit deficiencies that hinder their ability to effectively represent retrieval queries and documents. For instance, in the first example, although the anchor text “blood” is mentioned in the document, it is unrelated to the document’s theme.
We conducted statistical analysis on both the top 30 and bottom 30 groups. As displayed in Table 5, sentences from groups with high weights tend to be longer on average compared to those from groups with low weights. Since functional units (e.g. Learn More, Contact Us) are considered as separate sentences, we speculate that this preference is because functional units will hinder information extraction of the model. Furthermore, our analysis reveals that the model does not exhibit a clear preference between long and short queries. Additionally, the choice of frequent words indicates differing focuses between the top-weighted and bottom-weighted groups. The model appears to favor academic and journalistic content while marginalizing documents of an introductory or advertising nature.
Attributes & Features Top-Weighted Bottom-Weighted
30 Groups 30 Groups
Avg. Query Length 3.43 3.03
Avg. Sentence Length 9.34 8.20
Frequent Words
covid, news,
email, updated,
article, published,
health, secure,
data, privacy
encyc., new,
united, welcome,
university, help,
name, follow,
schools, contact
Table 5: Comparison between top-weighted and bottom-weighted groups. “Avg. Sentence Length” means the average number of words in a sentence in documents. “Frequent Words” are content words that most frequently appear in the corresponding groups.
6 Conclusion
This paper introduces WebDRO, an unsupervised dense retrieval model that utilizes web structures to cluster documents and adjusts group weights during contrastive training. We use structures of the web graph to train a link prediction style embedding model for clustering and utilize GroupDRO to reweight the clusters during training. Experiments on MS MARCO and BEIR show that WebDRO achieves significant improvements over baseline models, especially on difficult tasks. Further analysis shows that group weights learned by WebDRO are stable and valid, which means that the model’s preferences are consistent throughout training and that groups with larger weights are more beneficial for the model, while noisy or uninformative groups are down-weighted.
Limitations
In this work, we cluster different contrastive pairs by their semantic features, which means contrastive pairs with similar content will be clustered in the same group. This approach is based on a hypothesis: semantically resembled contrastive pairs have similar contributions to retrieval model training. However, other factors may also influence the data point’s contribution, like format, style, or even features we can’t explicitly describe. In future works, we hope to design a clustering strategy directly related to the optimization goal of dense retrieval to achieve more delicate clusters.
Research has shown that negative documents play a crucial part in dense retrieval, and introducing stronger (more confusing) negatives would enhance the model’s performance Xiong et al. (2021). In our methods, documents within each cluster are similar to each other, thus would serve as stronger negatives. However, to evaluate the impact of the reweighting strategy alone, we didn’t modify the way of choosing negative documents. In future works, we hope to make better use of our clusters and build stronger negatives to achieve better performance.
In this work, we mainly focus on the task of information retrieval itself. As information retrieval has vast applications in many fields, it would be interesting to see WebDRO’s performance in downstream applications like retrieval-augmented generation Guu et al. (2020) and retrieval-augmented diffusion Blattmann et al. (2022).
Appendix A Appendix
A.1 License
Our code base, OpenMatch, uses MIT License. For datasets used in this work, MS MARCO can be used freely without a license, BEIR uses Apache-2.0 license and ClueWeb22 uses an individual agreement, which we have already obtained.
A.2 Training Details
This subsection describes the training details of our WebDRO model, including data processing, hyperparameters and computational resources.
Data Processing. All training data are selected from the English documents in the B category of ClueWeb22. We follow Xie et al. (2023b) to process the training data by filtering out uninformative anchors (Homepage, Login, etc.), in-domain anchors, headers, and footers using heuristic rules. A miniBERT-based classifier model is also utilized to evaluate the quality of anchor texts. We only retain 25% anchor texts that are top-ranked.
Sizes of Datasets. The training data for the retrieval model contains $13.8$M contrastive pairs, and the training data for the embedding model contains $5.1$M contrastive pairs. For evaluation benchmarks, MS MARCO contains $8.8$M documents and $8,960$ queries, and BEIR contains $44.0$M documents and $47,282$ queries in total.
Hyperparameters. The learning rate and batch size are set to $53e-$ 5 and $384$ for training the embedding model. During the stage of training retrieval models, we set the learning rate of the contrastive loss to $53e-$ 5 and batch size to $768$. The learning rate for weight updates is set to $43e-$ 4.
Computational Resources. The training of our model uses 16 NVIDIA RTX A6000 GPUs (48GB of memory each). Each epoch of training (for all $13.8$M contrastive pairs) takes 9 to 12 hours depending on the inter-node communication efficiency.
A.3 More Details of Baseline Models
coCondenser Gao and Callan (2022) bases on the Condenser model Gao and Callan (2021) and is trained with token-level and span-level language modeling objectives. Contriever Izacard et al. (2021) constructs the positive pairs with random cropping and inverse cloze task. SPAR Chen et al. (2022) further trains dense retrievers on Wikipedia and aims at imitating a sparse retrieval teacher. AugTriever Meng et al. focuses on extracting the queries for the documents and using them to train retrieval models.
A.4 Analyse on the Number of Clusters
#Group MS MARCO Average on BEIR
50 26.93 40.40
200 27.43 40.48
500 27.37 40.70
1,000 27.11 40.49
5,000 26.82 40.17
15,000 27.29 40.00
Table 6: Retrieval performance of WebDRO with different group numbers (#Group). Results are in nDCG@10.
As shown in Sec. 3.2, we use K-Means to cluster the documents and the number of groups ($n$) is critical for GroupDRO. We conducted a series of experiments to investigate the impact of varying $n$ values on model performance. In Table 6, we can see the model performance initially is improved but later decreases as the $n$ increases. In our training scenario, $n=500$ is an appropriate value for WebDRO to achieve the best retrieval performance. If $n>500$, the number of the clustered groups is too large and the weights of smaller groups are usually unstable or even faulty updated. If $n<500$, the clustering is too coarse, and detailed features of documents cannot be represented through clustering. Each group contains a large number of documents, which should be assigned different weights.